# Optimizing the potential of KM3NeT in detecting core-collapse supernovae

### v1

@@ Abstract
Core-collapse supernovae mark the end of life of massive stars. However, despite their importance in astrophysics, their underlying mechanisms remain unclear. Neutrinos that emerge from the dense core of the star offer a promising way to study supernova dynamics. A strategy is presented to improve the potential of the KM3NeT neutrino telescope to detect core-collapse supernovae in our Galaxy or the Large Magellanic Cloud by further exploiting the properties of its optical modules equipped with multiple photomultipliers. A supernova burst is expected to produce a sudden hit rate increase in the KM3NeT detectors. New observables have been defined for individual optical modules that exploit the geometry and time distribution of the detected hits, enabling a better discrimination between signal and background signatures. In addition, a thorough investigation of the related systematic uncertainties is presented for the first time. When implemented, this new methodology allowed KM3NeT to probe 4646% more Galactic core-collapse supernova candidates than with the previous trigger strategy, reaching the dense Galactic bulge. It is now expected that, once completed, KM3NeT will achieve full Galactic sensitivity to core-collapse supernovae.
@@ 1 Introduction
Hot compact objects such as core-collapse supernovae (CCSNe) and binary neutron-star mergers are characterized by the thermal emission of MeV-scale neutrinos. The first and only observation of such neutrinos occurred in association with the SN1987A core-collapse supernova in the Large Magellanic Cloud [walker1987making]. This observation confirmed that around $99\%$ of the gravitational binding energy of a collapsing star is released in the form of neutrinos [schramm1990new], supporting the hypothesis that neutrino-driven mechanisms power supernova explosions. A total of only $24$ neutrinos were observed across the Kamiokande [hirata1988observation], IMB [bionta1987observation], and Baksan [alekseev1987possible] detectors, leaving significant gaps in our understanding. The mechanisms by which neutrino heating in the stalled shock wave leads to a successful explosion and asymmetric neutrino emission contributes to the explosion, remain subjects of active research [suzuki2024neutrinos]. To address these and other open questions [horiuchi2018can], detecting the next CCSN neutrino emission is essential. The CCSN neutrino signal precedes electromagnetic emissions from a supernova. Consequently, a fast combination of measurements from neutrino telescopes around the globe to triangulate the source’s position [al2021snews, coleiro2020combining] would allow for alerting optical telescopes to the upcoming visible counterpart. Such coordinated efforts between neutrino observatories and optical telescopes promise to provide a better understanding of these cataclysmic phenomena.
@@
In this paper, a new strategy is introduced, which improves the identification of low-energy signals in KM3NeT by characterizing the associated multi-hit coincidences using machine learning techniques. This strategy is used to provide an accurate estimate of low-energy backgrounds in KM3NeT, thus enhancing the detectability of CCSN neutrinos. A description of the KM3NeT detectors is given in Section 2. In Section 3, the the search for CCSN neutrinos with KM3NeT is presented and the set of new observables exploiting the structure of the KM3NeT DOMs is introduced. Section 4 describes how these new observables can be used to characterize the background signatures produced in KM3NeT by radioactive decays in the seawater and atmospheric muons. The new strategy based on machine learning to detect CCSN neutrino interactions is detailed in Section 5. The resulting sensitivity of KM3NeT to a CCSN neutrino burst is presented in Section 6. Finally, concluding remarks are given in Section 7.
@@ 2 The KM3NeT infrastructure
A significant challenge for a seawater detector like KM3NeT is the presence of large environmental backgrounds due to living organisms (bioluminescence) and ambient radioactivity. Characterizing and monitoring these backgrounds has been one of the main motivations for the multi-PMT design of the KM3NeT DOM [aiello2022km3net]. A DOM consists of $31$ three-inch photomultiplier tubes, enclosed in a pressure-resistant glass sphere of radius 22 $2\text{\,}\mathrm{cm}$. Each photomultiplier measures the time of detection of incoming photons with nanosecond accuracy.
@@
The data acquisition system of KM3NeT is based on the “all-data-to-shore” concept [KM3Net_LoI]. The data stream, consisting of all PMT hits above a threshold of $0.3$ photoelectrons, is subdivided into $100\text{\,}\mathrm{ms}$ segments called timeslices. The baseline hit rate within these timeslices is about $7\text{\,}\mathrm{kHz}$ per PMT, mainly due to radioactivity from 40K in water and to bioluminescence. Bioluminescence manifests itself as a superposition of sudden intense bursts of light above a slowly-varying seasonal component [priede2008potential]. To reduce the amount of these bursts, PMTs with a data-taking rate exceeding $20\text{\,}\mathrm{kHz}$ in a given timeslice are deactivated for the entire timeslice duration. This high-rate veto affects the performance of CCSN searches [aiello2022implementation]. In this article, following the convention from [aiello2022implementation], the impact of the high-rate-veto is assessed using the fraction of active PMTs, $f_{A}$, computed over all working DOMs. As in [van2021km3net], the capability of KM3NeT to detect CCSNe is evaluated for timeslices with $f_{A}>99$%. The impact of $f_{A}$ on the performance of KM3NeT is discussed in Section 6.
In the primary data stream, the hits stored in timeslices are used to identify triggered events, defined as groups of causally-connected hit coincidences distributed over multiple DOMs and corresponding to specific topologies. These triggers are used to identify and characterize high-energy neutrinos and atmospheric muons [KM3NeT:2024paj]. For CCSN neutrino studies, multi-hit coincidences within single DOMs are saved in a separate data stream called a SN timeslice. Such coincidences require at least 4 hits observed in a single DOM within a $15$ nanosecond window from PMTs whose axes lie within an angle of $90^{\circ}$[aiello2022implementation]. With these conditions, contributions from bioluminescence in the SN timeslice stream are negligible. Thus, with triggered events and SN timeslices, KM3NeT monitors neutrino interactions over two length scales: the inter-DOM scale, primarily sensitive to particles with energy in the GeV to PeV range, and the intra-DOM scale, which allows for probing particles down to an energy of a few MeVs.
@@ 3.1 Analysis overview
The analysis presented here aims at improving the performance of the search for CCSN neutrinos discussed in [van2021km3net] and [aiello2022implementation]. It adopts the same definition of a single-DOM event as a cluster of PMT hits observed in a DOM within a $10\text{\,}\mathrm{ns}$ time window. Single-DOM events are preselected by requiring that they belong to SN timeslices and pass a dedicated triggered event veto which removes events associated with highly energetic particles (e.g. atmospheric muons) and PMT afterpulses [van2021km3net].
In the CCSN analysis of [van2021km3net], the preselected single-DOM events were further filtered according to their multiplicity, i.e. the number of hits they contain. The analysis proposed here increases the KM3NeT sensitivity to CCSN neutrinos by considering, for each single-DOM event, not only the number of PMT hits but also their position, timing, and charge. Since a DOM contains $31$ PMTs, accounting for the properties of all the hits in each event would significantly increase the complexity of the analysis. To address this challenge, a set of single-DOM observables has been defined that reflects the properties of individual hits.
@@ 3.2 Single-DOM observables
  • The signal intensity is described by two observables: the multiplicity and the total time-over-threshold (ToT), defined as the sum of the time-over-thresholds of the PMT hits [ToT].
@@ 3.3 Comparison of signal and background
The multiplicity distributions for background and simulated signal events are shown in Figure 2. Intermediate multiplicities, between $6$ and $14$, are associated with the highest signal-to-background ratio. The shape of the background distribution, similar to a broken power law, reflects the background composition: radioactivity dominates up to multiplicity $7$ and atmospheric muons dominate above multiplicity $8$ [goos2023searching].
Figure 2: Multiplicity distributions for simulated CCSN signal events (orange) and background events from a $60$-day period of ORCA6 data (blue). The areas under the distributions are normalized. Only events passing the triggered event veto introduced in [van2021km3net] are included. For the signal, the $11~M_{\odot}$ progenitor from [tamborra2014neutrino] is used.
The distributions of total ToT, $|R|$, $\cos\theta$, and $\langle\Delta t\rangle$, for background and signal, are presented in Figure 3 for multiplicities $6$, 88, and $9$, corresponding to radioactivity-dominated, mixed, and muon-dominated backgrounds, respectively. For a fixed multiplicity, MeV-scale neutrinos, radioactive decays, and atmospheric muons produce distinct hit patterns. In all cases, a larger fraction of events with high-ToT is expected for the signal compared to the background. This can be explained by the fact that CCSN neutrino interactions close to DOMs can lead to large ToTs, while high-ToT background events are mostly associated with triggered muons, which are largely removed by the triggered event veto. Furthermore, the $|R|$, $\cos\theta$, and $\langle\Delta t\rangle$ distributions at low multiplicity (radioactivity-dominated region) and high multiplicity (muon-dominated region) show opposite trends, with the CCSN signal occupying an intermediate position (see also [goos2023searching]). This behavior relates to the fact that hits from radioactive decays are concentrated in both time and space since $\sim$$1\text{\,}\mathrm{MeV}$ radioactive decay products travel short distances and must be produced close to a DOM to be detected. Moreover, radioactive decay signatures are typically observed near the bottom of the DOM since the lower DOM hemisphere is the most densely instrumented. Conversely, hits from atmospheric muons are dispersed over a large area and time interval, as muons travel long distances, and are mostly observed in the upper DOM hemisphere since atmospheric muons are downward-going. Finally, CCSN signal events have topologies similar to radioactivity events but are associated with more energetic particles. Hence, they often have intermediate $|R|,\langle\Delta t\rangle,$ and $\cos\theta$ values and can thus be distinguished from both background components.
@@
Figure 3: Distributions of the total ToT (first row), hit concentration $|R|$ (second row), $\cos\theta$ (third row), and $\langle\Delta t\rangle$ (fourth row) for signal (orange) and background (blue) considering only events with multiplicity 66 (left), 88 (center), and 99 (right). The simulated signal is based on the $11M_{\odot}$ model from [tamborra2014neutrino]. The background is taken from ORCA6 data and is composed of atmospheric muons and radioactive decays. The areas under all the distributions are normalized to unity. Only events that passed the triggered event veto are included.
@@ 4 Background characterization at KM3NeT
In this section, a data-driven method is presented to estimate the radioactivity and muon background rates. In the 661010 multiplicity range used in CCSN studies, radioactivity and muon backgrounds coexist and their relative proportions have, to date, been evaluated using simulations [radioactivity_simulation]. Measuring the muon and radioactivity rates separately for each multiplicity is essential to estimate the sensitivity of upcoming ARCA and ORCA configurations to CCSN neutrinos and assess the accuracy of the KM3NeT detector simulations. Beyond the CCSN analysis, estimating the radioactivity background rates over a wide range of multiplicities could enhance searches for other MeV-scale phenomena related to e.g. beyond-the-Standard-Model physics, and improve the modeling of high-energy neutrino signatures in ARCA and ORCA.
Although the total background rates in individual DOMs can be directly measured for all multiplicities [dom_rates], determining the relative fractions of atmospheric muon and radioactivity events for a given multiplicity can be challenging. This is especially true in the CCSN multiplicity search region, where both backgrounds are present in comparable amounts once the event preselection described in Section 3 is applied. However, radioactivity and atmospheric muons leave distinctive hit patterns on individual DOMs, which can be distinguished using the single-DOM observables introduced in Section 3.
@@ 4.2 Estimating radioactivity and muon background rates
For all tested samples, the muon rates do not vary significantly with the multiplicity and remain within $15\%$ of the simulation predictions. Conversely, the radioactivity event rates evaluated using this data-driven method for multiplicities $5$ to $10$ differ by up to $60\%$ from the simulated rates, as shown in Figure 5. These rates also significantly differ from the values used in the previous CCSN analysis [van2021km3net], which were exclusively based on previous, simplified, simulation tools. Although these differences are expected to have a negligible impact on the results from [van2021km3net], they could sizably affect multivariate analyses, such as the one presented in Sections 5 and 6, where this method is used to evaluate systematic uncertainties on the CCSN signal acceptance (i.e. the fraction of simulated CCSN events that survive the selection), and to estimate background levels for the final ARCA and ORCA detector configurations. Improving low-energy background estimates using single-DOM observables will allow for refining reconstruction algorithms and reducing systematic uncertainties across various KM3NeT searches.
Figure 5: Radioactivity rates predicted by the simulations described in Appendix B.2.1 (dashed line) and rescaled for each multiplicity using the procedure described in Section 4.2 (blue triangles). The uncertainty bars on the rescaled rates account for both the finite statistics in the control regions and the modeling uncertainties on the background fractions. The red dots indicate the radioactivity rate estimations obtained using simulations tools, as in the previous CCSN analysis [van2021km3net].
@@ 5.1 Identifying CCSN signals and selecting observables using Boosted Decision Trees
Boosted Decision Trees (BDTs) are trained to separate CCSN neutrino signatures from the background events that survive the triggered event veto. The signal is simulated assuming the $11~M_{\odot}$ model of [tamborra2014neutrino]; the choice of supernova model has been found to have little impact on the performance. Data from the ORCA6 and ARCA21 detector configurations are used for modeling the background. Since the background composition in the multiplicity range 595-9 strongly depends on both the detector structure and the multiplicity, five BDTs are used for ARCA and ORCA separately: one for each multiplicity from $5$ to $8$ and one for multiplicities $\geq 9$.
The BDTs employed for this analysis are built with the gradient boosting method of the XGBoost library [Chen:2016:XST:2939672.2939785]. For each multiplicity and detector configuration, the BDTs are trained using various subsets of the observables introduced in Section 3.2 as features, and the classification accuracy is optimized using a mean-squared-error loss function. The predictability of the BDT is ensured through cross-validation, with a 33:11 splitting between the training and test samples. Finally, the BDT hyperparameters are optimized by maximizing the background rejection for a benchmark signal acceptance of 50%.

### v2

@@ Abstract
Core-collapse supernovae mark the end of life of massive stars. However, despite their importance in astrophysics, their underlying mechanisms remain unclear. Neutrinos that emerge from the dense core of the star offer a promising way to study supernova dynamics. A strategy is presented to improve the potential of the KM3NeT neutrino telescope to detect core-collapse supernovae in our Galaxy or the Large Magellanic Cloud by further exploiting the properties of its optical modules equipped with multiple photomultipliers. A supernova burst is expected to produce a sudden hit rate increase in the KM3NeT detectors, which could be used to detect a supernova even in the absence of triggers from other experiments. New observables have been defined for individual optical modules that exploit the geometry and time distribution of the detected hits, enabling a better discrimination between signal and background signatures. In addition, a thorough investigation of the related systematic uncertainties is presented for the first time. When implemented, this new methodology allowed KM3NeT to probe 4646% more Galactic core-collapse supernova candidates than with the previous trigger strategy, reaching the dense Galactic bulge. It is now expected that, once completed, KM3NeT will achieve full Galactic sensitivity to core-collapse supernovae independently from other experiments.
@@ 1 Introduction
Hot compact objects such as core-collapse supernovae (CCSNe) and binary neutron-star mergers are characterized by the thermal emission of MeV-scale neutrinos. The first and only observation of such neutrinos occurred in association with the SN1987A core-collapse supernova in the Large Magellanic Cloud [30]. This observation confirmed that around $99\%$ of the gravitational binding energy of a collapsing star is released in the form of neutrinos [25], supporting the hypothesis that neutrino-driven mechanisms power supernova explosions. A total of only about $24$ neutrinos were observed across the Kamiokande [21], IMB [12], and Baksan [10] detectors, leaving significant gaps in our understanding. The mechanisms by which neutrino heating in the stalled shock wave leads to a successful explosion and asymmetric neutrino emission contributes to the explosion, remain subjects of active research [26]. To address these and other open questions [22], detecting the next CCSN neutrino emission is essential. The CCSN neutrino signal precedes electromagnetic emissions from a supernova. Consequently, a fast combination of measurements from neutrino telescopes around the globe to triangulate the source’s position [8, 16] would allow for alerting optical telescopes to the upcoming visible counterpart. Such coordinated efforts between neutrino observatories and optical telescopes promise to provide a better understanding of these cataclysmic phenomena.
@@
In this paper, a new strategy is introduced, which improves the identification of low-energy signals in KM3NeT by characterizing the associated multi-hit coincidences using machine learning techniques. This strategy is used to provide an accurate estimate of low-energy backgrounds in KM3NeT, thus enhancing the detectability of CCSN neutrinos. A description of the KM3NeT detectors is given in Section 2. In Section 3, the search for CCSN neutrinos with KM3NeT is presented and the set of new observables exploiting the structure of the KM3NeT DOMs is introduced. Section 4 describes how these new observables can be used to characterize the background signatures produced in KM3NeT by radioactive decays in the seawater and atmospheric muons. The new strategy based on machine learning to detect CCSN neutrino interactions is detailed in Section 5. The resulting sensitivity of KM3NeT to a CCSN neutrino burst is presented in Section 6. Finally, concluding remarks are given in Section 7.
@@ 2 The KM3NeT infrastructure
A significant challenge for a seawater detector like KM3NeT is the presence of large environmental backgrounds due to living organisms (bioluminescence) and ambient radioactivity. Characterizing and monitoring these backgrounds has been one of the main motivations for the multi-PMT design of the KM3NeT DOM [7]. A DOM consists of $31$ three-inch photomultiplier tubes, enclosed in a pressure-resistant glass sphere of radius 22 $2\text{\,}\mathrm{cm}$. Each photomultiplier measures the time of detection of incoming photons with nanosecond accuracy.
@@
The data acquisition system of KM3NeT is based on the “all-data-to-shore” concept [3]. The data stream, consisting of all PMT hits above a threshold of $0.3$ photoelectrons, is subdivided into $100\text{\,}\mathrm{ms}$ segments called timeslices. The baseline hit rate within these timeslices is about $7\text{\,}\mathrm{kHz}$ per PMT, mainly due to radioactivity from 40K in water and to bioluminescence. Bioluminescence manifests itself as a superposition of sudden intense bursts of light above a slowly-varying seasonal component [23]. To reduce the amount of these bursts, PMTs with a data-taking rate exceeding $20\text{\,}\mathrm{kHz}$ in a given timeslice are deactivated for the entire timeslice duration. This high-rate veto affects the performance of CCSN searches [6]. In this article, following the convention from [6], the impact of the high-rate-veto is assessed using the fraction of active PMTs, $f_{A}$, computed over all working DOMs. As in [5], the capability of KM3NeT to detect CCSNe is evaluated for timeslices with $f_{A}>99$%. The impact of $f_{A}$ on the performance of KM3NeT is discussed in Section 6.
In the primary data stream, the hits stored in timeslices are used to identify triggered events, defined as groups of causally-connected hit coincidences distributed over multiple DOMs and corresponding to specific topologies. These triggers are used to identify and characterize high-energy neutrinos and atmospheric muons [4]. For CCSN neutrino studies, multi-hit coincidences within single DOMs are saved in a separate data stream called a SN timeslice. Such coincidences require at least 4 hits observed in a single DOM within a $15$ nanosecond window from PMTs whose axes lie within an angle of $90^{\circ}$[6]. With these conditions, contributions from bioluminescence in the SN timeslice stream are negligible. Thus, with triggered events and SN timeslices, KM3NeT monitors neutrino interactions over two length scales: the inter-DOM scale, primarily sensitive to particles with energy in the GeV to PeV range, and the intra-DOM scale, which allows for probing particles down to an energy of a few MeVs.
@@ 3.1 Analysis overview
The analysis presented here aims at improving the performance of the search for CCSN neutrinos discussed in [5] and [6]. It adopts the same definition of a single-DOM event as a cluster of PMT hits observed in a DOM within a $10\text{\,}\mathrm{ns}$ time window. Single-DOM events are preselected by requiring that they belong to SN timeslices and pass a dedicated triggered event veto which removes events associated with highly energetic particles (e.g. atmospheric muons) and PMT afterpulses [5].
In the CCSN analysis of [5], the preselected single-DOM events were further filtered according to their multiplicity, i.e. the number of hits they contain. The analysis proposed here increases the KM3NeT sensitivity to CCSN neutrinos by considering, for each single-DOM event, not only the number of PMT hits but also their position, timing, and charge. Since a DOM contains $31$ PMTs, accounting for the properties of all the hits in each event would significantly increase the complexity of the analysis. To address this challenge, a set of single-DOM observables has been defined that reflects the properties of individual hits.
@@ 3.2 Single-DOM observables
  • The signal intensity is described by two observables: the multiplicity and the total time-over-threshold (ToT), defined as the sum of the time-over-thresholds of the PMT hits [13].
@@ 3.3 Comparison of signal and background
The multiplicity distributions for background and simulated signal events are shown in Figure 2. Intermediate multiplicities, between $6$ and $14$, are associated with the highest signal-to-background ratio. The shape of the background distribution, similar to a broken power law, reflects the background composition: radioactivity dominates up to multiplicity $7$ and atmospheric muons dominate above multiplicity $8$ [19].
Figure 2: Multiplicity distributions for simulated CCSN signal events (orange) and background events from a $60$-day period of ORCA6 data (blue). The areas under the distributions are normalized. Only events passing the triggered event veto introduced in [5] are included. For the signal, the $11~M_{\odot}$ progenitor from [27] is used.
The distributions of total ToT, $|R|$, $\cos\theta$, and $\langle\Delta t\rangle$, for background and signal, are presented in Figure 3 for multiplicities $6$, 88, and $9$, corresponding to radioactivity-dominated, mixed, and muon-dominated backgrounds, respectively. For a fixed multiplicity, MeV-scale neutrinos, radioactive decays, and atmospheric muons produce distinct hit patterns. In all cases, a larger fraction of events with high-ToT is expected for the signal compared to the background. This can be explained by the fact that CCSN neutrino interactions close to DOMs can lead to large ToTs, while high-ToT background events are mostly associated with triggered muons, which are largely removed by the triggered event veto. Furthermore, the $|R|$, $\cos\theta$, and $\langle\Delta t\rangle$ distributions at low multiplicity (radioactivity-dominated region) and high multiplicity (muon-dominated region) show opposite trends, with the CCSN signal occupying an intermediate position (see also [19]). This behavior relates to the fact that hits from radioactive decays are concentrated in both time and space since $\sim$$1\text{\,}\mathrm{MeV}$ radioactive decay products travel short distances and must be produced close to a DOM to be detected. Moreover, radioactive decay signatures are typically observed near the bottom of the DOM since the lower DOM hemisphere is the most densely instrumented. Conversely, hits from atmospheric muons are dispersed over a large area and time interval, as muons travel long distances, and are mostly observed in the upper DOM hemisphere since atmospheric muons are downward-going. Finally, CCSN signal events have topologies similar to radioactivity events but are associated with more energetic particles. Hence, they often have intermediate $|R|,\langle\Delta t\rangle,$ and $\cos\theta$ values and can thus be distinguished from both background components.
@@
Figure 3: Distributions of the total ToT (first row), hit concentration $|R|$ (second row), $\cos\theta$ (third row), and $\langle\Delta t\rangle$ (fourth row) for signal (orange) and background (blue) considering only events with multiplicity 66 (left), 88 (center), and 99 (right). The simulated signal is based on the $11M_{\odot}$ model from [27]. The background is taken from ORCA6 data and is composed of atmospheric muons and radioactive decays. The areas under all the distributions are normalized to unity. Only events that passed the triggered event veto are included.
@@ 4 Background characterization at KM3NeT
In this section, a data-driven method is presented to estimate the radioactivity and muon background rates. In the 661010 multiplicity range used in CCSN studies, radioactivity and muon backgrounds coexist and their relative proportions have, to date, been evaluated using simulations [20]. Measuring the muon and radioactivity rates separately for each multiplicity is essential to estimate the sensitivity of upcoming ARCA and ORCA configurations to CCSN neutrinos and assess the accuracy of the KM3NeT detector simulations. Beyond the CCSN analysis, estimating the radioactivity background rates over a wide range of multiplicities could enhance searches for other MeV-scale phenomena related to e.g. beyond-the-Standard-Model physics, and improve the modeling of high-energy neutrino signatures in ARCA and ORCA.
Although the total background rates in individual DOMs can be directly measured for all multiplicities [2], determining the relative fractions of atmospheric muon and radioactivity events for a given multiplicity can be challenging. This is especially true in the CCSN multiplicity search region, where both backgrounds are present in comparable amounts once the event preselection described in Section 3 is applied. However, radioactivity and atmospheric muons leave distinctive hit patterns on individual DOMs, which can be distinguished using the single-DOM observables introduced in Section 3.
@@ 4.2 Estimating radioactivity and muon background rates
For all tested samples, the muon rates do not vary significantly with the multiplicity and remain within $15\%$ of the simulation predictions. Conversely, the radioactivity event rates evaluated using this data-driven method for multiplicities $5$ to $10$ differ by up to $60\%$ from the simulated rates, as shown in Figure 5. These rates also significantly differ from the values used in the previous CCSN analysis [5], which were exclusively based on previous, simplified, simulation tools. Although these differences are expected to have a negligible impact on the results from [5], they could sizably affect multivariate analyses, such as the one presented in Sections 5 and 6, where this method is used to evaluate systematic uncertainties on the CCSN signal acceptance (i.e. the fraction of simulated CCSN events that survive the selection), and to estimate background levels for the final ARCA and ORCA detector configurations. Improving low-energy background estimates using single-DOM observables will allow for refining reconstruction algorithms and reducing systematic uncertainties across various KM3NeT searches.
Figure 5: Radioactivity rates predicted by the simulations described in Appendix B.2.1 (dashed line) and rescaled for each multiplicity using the procedure described in Section 4.2 (blue triangles). The uncertainty bars on the rescaled rates account for both the finite statistics in the control regions and the modeling uncertainties on the background fractions. The red dots indicate the radioactivity rate estimations obtained using simulations tools, as in the previous CCSN analysis [5].
@@ 5.1 Identifying CCSN signals and selecting observables using Boosted Decision Trees
Boosted Decision Trees (BDTs) are trained to separate CCSN neutrino signatures from the background events that survive the triggered event veto. The signal is simulated assuming the $11~M_{\odot}$ model of [27]; the choice of supernova model has been found to have little impact on the performance. Data from the ORCA6 and ARCA21 detector configurations are used for modeling the background. Since the background composition in the multiplicity range 595-9 strongly depends on both the detector structure and the multiplicity, five BDTs are used for ARCA and ORCA separately: one for each multiplicity from $5$ to $8$ and one for multiplicities $\geq 9$.
The BDTs employed for this analysis are built with the gradient boosting method of the XGBoost library [15]. For each multiplicity and detector configuration, the BDTs are trained using various subsets of the observables introduced in Section 3.2 as features, and the classification accuracy is optimized using a mean-squared-error loss function. The predictability of the BDT is ensured through cross-validation, with a 33:11 splitting between the training and test samples. Finally, the BDT hyperparameters are optimized by maximizing the background rejection for a benchmark signal acceptance of 50%.